# Relationship Between Major Stellar Physical Parameters and Normal Mode Frequencies in Accreting White Dwarf Stars

### v1

@@ Acknowledgements
This manuscript has benefited from insightful comments by the anonymous referee. We thank Ken Shen, Alan Calder, Sam Boos, and Nethra Rajavel for the useful discussions. We also would like to thank Spencer Caldwell and Broxten Miles for their earlier work on setting up the nova simulation. This work was supported under programs HST-GO-15072, HST-GO-16069, and HST-AR-16638 through the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5-26555. Support for these programs was provided through a grant from the STScI under NASA contract NAS5-26555.
Software: MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019, https://docs.mesastar.org/en/latest/), GYRE (Townsend & Teitler, 2013; Townsend et al., 2018, https://gyre.readthedocs.io/en/stable/) Matplotlib (Hunter, 2007)

### v2

@@ Acknowledgements
This manuscript has benefited from insightful comments by the anonymous referee. We thank Ken Shen, Alan Calder, Sam Boos, and Nethra Rajavel for the useful discussions. We also would like to thank Spencer Caldwell and Broxton Miles for their earlier work on setting up the nova simulation. This work was supported under programs HST-GO-15072, HST-GO-16069, and HST-AR-16638 through the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5-26555. Support for these programs was provided through a grant from the STScI under NASA contract NAS5-26555.
Software: MESA (Paxton et al. 2011, 2013, 2015, 2018, 2019, https://docs.mesastar.org/en/latest/),
GYRE (Townsend & Teitler 2013; Townsend et al. 2018, https://gyre.readthedocs.io/en/stable/),
Matplotlib (Hunter 2007, https://matplotlib.org/)
ORCID iDs
Praphull Kumar: 0000-0002-8791-3704
Dean M. Townsley: 0000-0002-9538-5948